OpenStudy (anonymous):

x+y=62.95 35x + 20y =200

OpenStudy (anonymous):

Do elimination method. Multiply the first equation so that you can eliminate "y" by subtracting the two equations.

OpenStudy (anonymous):

x + y = 62.95 ------- (1) 35x + 20y = 200 ---- (2) Multiply (1) by 20. You get: 20x + 20y = 20*62.95 35x + 20y = 200 Subtracting, you get: -15x = 20*62.95 - 200 x = -(20*62.95-200)/15 Find y using (1) above.